How much joules of heat will it take to increase the temperature of 6kg water 5 k ?


1
Expert's answer
2021-01-12T12:12:19-0500

We can find the amount of heat from the formula:


"Q=mc\Delta T,""Q=6\ kg\cdot 4180\ \dfrac{J}{kg\cdot K}\cdot 5\ K=1.254\cdot10^5\ J."

Answer:

"Q=1.254\cdot10^5\ J."